# Quasiconservation Laws and Suppressed Transport in Weakly Interacting Localized Models

### v2

@@ III.4.1 LIOM Corrections: Anderson Model
Just as in the distributions of $\|\hat{X}\|_{F}$, the long $\|\hat{X}\|_{F}^{-2}$ power-law tails appear here for the same reason as they did in the AGP distributions [see the inset in Fig. 3(d)]. This shows that the perturbation theory breaks down for a fraction of the LIOMs, which receive large corrections due to resonances. These resonances are well-known in studies of MBL, where different methods have been developed to treat them. One example is provided in Ref. [76]. In that work, the authors separate the Hamiltonian into resonating and non-resonating regions, where resonances involve small denominators past some cutoff $\delta$. They then treat these resonating regions nonperturbatively, which is possible given the small sizes of those resonating regions. In our work, we do not attempt to deliver any special treatment to these resonances since our work is, in spirit, perturbative. Rather, our goal is to extract as much insight as possible from the perturbative expansion, while being aware of the presence of resonances and accounting for their effects in the naive perturbative treatment.
III.4.2 LIOM Corrections: Quasiperiodic Model
@@
Figure 5: Analysis of the Frobenius norm of the first-order corrections to the noninteracting LIOMs, $\|\tilde{n}_{\alpha}^{(1)}\|_{F}$, for the weakly-interacting quasiperiodic model. (a)-(d): Distributions of $\ln(\|\tilde{n}_{\alpha}^{(1)}\|_{F})$ based on 500 phase shifts $\delta$, sampled at potential depths $h=0.6,1.6,2.6,3.1$, with $\alpha=1\dots L$. Each curve represents the distribution calculated at different system sizes, from $L=20$ to $L=100$ in steps of 10. The inset in (d) shows the same data as in the main (d) panel but with the vertical scale shown on a log scale. The red line in the inset represents a polynomial $\|\tilde{n}_{\alpha}^{(1)}\|_{F}^{-1.75}$ fit to the tail of the probability distribution of $\|\tilde{n}_{\alpha}^{(1)}\|_{F}$— see the text surrounding Eq. (21) for details on the form plotted. Similarly to what was observed for the interacting Anderson model, the quasiperiodic model seems to admit an extensive number of quasiconservation laws for the localized phase $h>2$. (e): Median $\mathrm{med}\|\tilde{n}_{\alpha}^{(1)}\|_{F}$ as a function of $L$ on a log-log scale. Each colored curve represents a different potential depth $h$, from $h=0.1$ to $h=3.1$ in steps of 0.5. One observes that the curves for $h<2$ grow with $L$, indicating that the corrections are, unsurprisingly, delocalized. For $h>2$ the curves saturate, indicating that quasi-LIOMs persist when interactions are added to the localized regime of the quasiperiodic model. For a full discussion of the results, see Sec. III.4.2
@@
III.4.3 LIOM Corrections: Conclusions
Our numerical results show that, even when perturbed by weak interactions, quasi-LIOMs continue to exist: specifically, we saw that the two localized models considered still preserve an $O(1)$ fraction of the noninteracting LIOMs up to first order. These LIOMs hence survive until time scales of $O(\lambda^{-4})$, supporting that thermalization happens on anomalously slow timescales in these systems. From the perspective of integrability breaking, this makes interactions a “weak” perturbation to localized models– consistent with the existence of MBL. However, the results raise many questions. In particular, the slow tails in the distributions of $\|\tilde{n}^{(1)}_{\alpha}\|_{F}$ show that perturbation theory breaks down for a finite fraction of the LIOMs— the fate of those LIOMs remains unclear when interactions are added. From just the norms of the corrections, it is unclear what the spatial structures of the resonances are that cause the breakdown of these LIOMs—in particular, do the resonances mix LIOMs locally, or nonlocally? In principle, some such locality information is contained in the operator content of the corresponding corrections, but it is not easy to analyze. In the next section, we introduce the charge transport capacity, which can frame this question from a different and very physically relevant perspective.
@@ IV.1 Charge Transport Capacity
The charge transport capacity over a time $t$ is defined as the operator norm of
[…]
@@
In the MBL phase, only LIOMs localized near $i_{0}$ are expected to contribute appreciably to the charge transport capacity, which should remain bounded and not scale with the system size.
To make the charge transport capacity practically computable and to quantify this intuition for the MBL phase, we note that the relation
@@
To tightly bound the charge transport capacity, we aim to find $\hat{D}$ such that the operator norm of $\hat{M}_{i_{0}}$ is minimized. If such $\hat{M}_{i_{0}}$ is localized, its norm is finite in the thermodynamic limit, and the charge transport capacity remains bounded for arbitrarily long times. Since this quantity is defined as the operator norm of $\Delta\hat{N}_{i_{0}}$, this would prove the absence of transport for any initial state to infinite times. In practice, minimizing the operator norm of $\hat{M}_{i_{0}}$ may be challenging, and it is typically easier to minimize its Frobenius norm. Nevertheless, we expect that if the system is localized, the resulting $\hat{M}_{i_{0}}$ will be sufficiently quasilocal, so that the Frobenius norm serves as a good proxy for the operator norm, up to an overall numerical factor. Indeed, in a localized system, only degrees of freedom near the link $i_{0}$ would contribute significantly to the transport capacity, and thus $\hat{M}_{i_{0}}$ should be quasilocal.
So far, we have only presented a heuristic discussion of the behavior of $\|\hat{M}_{i_{0}}\|_{\text{op}}$. Below, we proceed with some general comments on the properties of $\|\hat{M}_{i_{0}}\|_{\text{op}}$, as well as a more formal argument that $\|\hat{M}_{i_{0}}\|_{\text{op}}$ is finite in the thermodynamic limit if the Hamiltonian can be written in terms of LIOMs (i.e., as in Eq. (7)). Denoting many-body eigenstates/energies of $H$ as {|$\{\ket{\mu},E_{\mu}\}$, and assuming non-degenerate spectrum (which is reasonable for a finite-size disordered system), the matrix elements of the many-body operator $\hat{M}_{i_{0}}$ for $\mu\neq\nu$ are uniquely fixed
$\big{(}\hat{M}_{i_{0}}\big{)}_{\mu\nu}=\frac{(\hat{J}_{i_{0}})_{\mu\nu}}{i(E_{\mu}-E_{\nu})}=\big{(}\hat{N}_{R}\big{)}_{\mu\nu}\leavevmode\nobreak\ .$ (26)
@@
$\hat{M}_{i_{0}}=[\hat{N}_{R}(i_{0})]_{\text{off-diag}}=\hat{N}_{R}-[\hat{N}_{R}]_{\text{diag}}\leavevmode\nobreak\ .$ (27)
@@
From now on, we will consider $\hat{M}_{i_{0}}$ given by Eq. (27) and we will call this operator the charge transport capacity. Note that this operator is easy to calculate numerically if one knows the full spectrum of $\hat{H}$; this could be an interesting new measure to explore in the full interacting problem, although in this work we will primarily focus on perturbative characterizations.
@@
We note that the above argument relies on the validity of the LIOM model in the MBL phase. However, the possible double-logarithmic growth of the number entropy in the MBL regime, as reported in Ref. [77], suggests that more subtle transport phenomena may persist even deep within the localized phase; whether this is true is not entirely resolved in the literature [78, 79]. Regardless, we emphasize that our calculations of the charge transport in this paper are fully independent of any assumptions about the form of the LIOM description in the localized regime.
@@ IV.2 Analytical Forms of Noninteracting Transport and Its Perturbative Correction
We thus find that $\hat{M}^{(1)}$ has to satisfy the relation $[\hat{H}_{0},\hat{M}^{(1)}]+[\hat{V},\hat{M}^{(0)}]=0$. In this case, $\hat{M}^{(1)}$ is a four-fermion operator:
@@ IV.3.1 Transport in the Single-Particle Anderson Model
Figure 9: Analysis of $\|\hat{M}^{(0)}\|_{F}$ for the noninteracting Anderson Model based on 10000 disorder realizations, open boundary conditions, and with $i_{0}=\lfloor L/2\rfloor$. (a)–(d): Calculated probability density functions of $\|\hat{M}^{(0)}\|_{F}$. Each panel represents different disorder strengths $W=0.6,1.6,2.6,3.1$. Within each panel, each colored curve represents different system sizes from $L=20$ to $L=100$ in steps of 101010. The distributions are converged at large $L$, consistent with the presence of localization for arbitrarily weak disorder. (e): Medians of the distributions in (a)-(d), $\mathrm{med}\|\hat{M}^{(0)}\|_{F}$, as a function of $L$. Each curve represents a different disorder strength. Here, one sees that the curves saturate at large $L$, reflecting that, in the thermodynamic limit, charge cannot freely flow in the single-particle Anderson model due to localization. In the no disorder case $W=0$, the median scales as $\approx\frac{1}{4}\sqrt{L}$.
@@ IV.3.2 Transport in the Single-Particle Quasiperiodic Model
Figure 11: Analysis of the Frobenius norm of the charge transport capacity, $\|\hat{M}^{(0)}\|_{F}$, for the noninteracting quasiperiodic model with open boundary conditions. The analysis is done based on 10000 evenly spaced phase shifts $\delta$ between $\delta=0$ and $\delta=\frac{1}{k}$, with $i_{0}=\lfloor L/2\rfloor$, and wavenumber $k=\sqrt{2}$. (a)–(d): Calculated probability density functions of $\|\hat{M}^{(0)}\|_{F}$. Within each panel, each colored curve represents a different system size, going from $L=20$ to $L=100$, in steps of 101010. Panels (a) and (b) are in the delocalized phase for $h=0.6$ and 1.61.61.6, where one sees narrow and rightwards shifting peaks. This is a signature of unbounded transport in the delocalized phase. The localized phase is represented in panels (c) and (d) for potential depths $h=2.6,3.1$, which shows distributions converging sharply with increasing $L$. This reflects that transport is finite and limited in the localized phase. (e): We show the median of $\|\hat{M}^{(0)}\|_{F}$, $\mathrm{med}\|\hat{M}^{(0)}\|_{F}$, as a function of $L$. Each differently colored curve represents a different potential depth $h$, from $h=0.1$ to $h=3.1$ in steps of 0.5. We also show the curve at the transition $h=2$ (yellow). The curves in the extended phase $h<2$ show a square root scaling similar to that in the no-disorder case, while the curves in the localized phase for $h>2$ saturate to finite values at large $L$.
@@ IV.4.1 Anderson Model
In Fig. 12, we show numerical results for $\|\hat{M}^{(1)}\|_{F}$ in the weakly-interacting Anderson model.
@@
Figure 12: Analysis of the Frobenius norm of the first-order correction to the charge transport capacity, $\|\hat{M}^{(1)}\|_{F}$, for the weakly-interacting Anderson Model across the link with $i_{0}=\lfloor L/2\rfloor$ and open boundary conditions. (a)–(d): Calculated probability density functions of $\ln(\|\hat{M}^{(1)}\|_{F})$ based on 1000 disorder realizations. Each panel shows a different disorder strength $W=0.6,1.6,2.6,3.1$. Within each panel, each colored curve represents a different system size, with system sizes from $L=20$ to $L=100$ in steps of 10. The inset in (d) shows a distribution of $\|\hat{M}^{(1)}\|_{F}$ at $h=3.1$ with a natural log scale on the y and x-axes, with a $\|\hat{M}^{(1)}\|_{F}^{-2}$ (red line) comparison to the tail— see the text surrounding Eq. (21) for details on the form plotted. The convergence of the distributions at large disorder and $L$ indicates that the corrections to transport are bounded in the thermodynamic limit, suggesting that charge transport is suppressed in the presence of interactions at first-order in perturbation theory. (e): Plot of $\mathrm{med}\|\hat{M}^{(1)}\|_{F}$ as a function of $L$. Each curve represents a different disorder strength $W$, from $W=0.6$ to 3.13.13.1 in steps of 0.5. The inset in (e) shows the same data but plotted using a log-log scale. At large $L$, the transport is bounded, indicating the presence of bottlenecks even when interactions are added.
@@
To check this, we numerically calculate whether or not transport corrections larger than the median at a site $i_{0}$ will continue to stay larger than the median as one travels away from $i_{0}$. Thus, we define $\Theta_{i}\equiv\Theta\left(\|\hat{M}_{i}^{(1)}\|_{F}-\mathrm{med}\|\hat{M}_{i}^{(1)}\|_{F}\right)$. Here, $\Theta$ refers to the Heaviside function—thus, $\Theta_{i}$ is 1 if $\|\hat{M}_{i}^{(1)}\|_{F}$ is larger than the median of the $\|\hat{M}_{i}^{(1)}\|_{F}$ distribution, and 0 otherwise. We calculate the correlation coefficient $G(D)$ between $\Theta_{L/2}$ and $\Theta_{L/2+D}$ (for concision, we denote $\lfloor L/2\rfloor=L/2$, which is true in this case since we only perform calculations on lattices of even sizes), and limit ourselves to considering cases, out of 500500500 disorder realizations, where $\|\hat{M}_{L/2}^{(1)}\|_{F}$ are large charge transport corrections (defined as corrections greater than the median of $\|\hat{M}_{L/2}^{(1)}\|_{F}$ for those 500 disorder realizations) at the central link $i_{0}=\lfloor L/2\rfloor$. The full expression for this correlation function is

### v3

@@ III.4.1 LIOM Corrections: Anderson Model
Resonances between noninteracting orbitals (not to be confused with many-body resonances) are known to play a big role in MBL, where different methods have been developed to treat them. One example is provided in Ref. [78]. In that work, the authors separate the Hamiltonian into resonating and non-resonating regions, where resonances involve small denominators past some cutoff $\delta$. They then treat these resonating regions nonperturbatively, which is possible given the small sizes of those resonating regions. In our work, we do not attempt to deliver any special treatment to these resonances since our work is, in spirit, perturbative. Rather, our goal is to extract as much insight as possible from the perturbative expansion, while being aware of the presence of resonances and accounting for their effects in the naive perturbative treatment.
III.4.2 Dependence of LIOM Corrections on Energy and Participation Ratio
Figure 5: A density scatter plot of the natural log of the LIOM corrections $\ln\|\tilde{n}^{(1)}_{\alpha}\|_{F}$ at $W=3.1$ and $L=80$ against (a) the energy $E$ of its corresponding orbital $\alpha$ and (b) the participation ratio (PR) of the same orbital. Here, we use 500 different disorder realizations. For each disorder realization, we keep corrections to all orbitals $\alpha=1,\dots,L$. Yellower colors indicate a higher density of points.
The defining characteristic of the distributions of $\|\tilde{n}^{(1)}_{\alpha}\|_{F}$ is the presence of slowly decaying power-law tails, which cause the mean to diverge. These tails also occur with large disorder strengths. To investigate the origin of these large corrections, we compute the dependence of the LIOM corrections on both the orbital energy and participation ratio when the disorder is larger at $W=3.1$, and show our results in Fig. 5. One might intuitively expect that large corrections arise from rare thermal regions where the variation in the onsite potential is anomalously small. However, the data in Fig. 5 tells a more complex and partly counterintuitive story.
Figure 5(a) shows that the typical LIOM corrections are roughly correlated with energy, with larger corrections in the middle of the spectrum, although the correlation is not strong as long as one is away from the edges of the spectrum. This is what one would expect, as states near the middle of the spectrum tend to be more delocalized. Delocalized orbitals typically produce larger numerators in the expression for the matrix elements of the AGP, Eq. (20). This trend is echoed in panel (b), where the typical and lower-bound LIOM corrections increase with the participation ratio (PR) of the corresponding orbital.
In contrast, the very large LIOM corrections that form the long tails tell a counterintuitive story. In Fig. 5(b), these large corrections appear anticorrelated with the participation ratio: they mostly occur with orbitals with a small PR, and seem less likely for those with large PR. A possible explanation is that rare thermal regions with large participation ratios exhibit stronger level repulsion, which suppresses small denominators in Eq. (20). With this, it is reasonable to conjecture that the dominant mechanism behind the largest LIOM corrections is not rare thermal regions, but rather energy resonances that produce small four-energy denominators in Eq. (20). Since the numerators would be very small for orbitals that are spatially distant, we can further guess that these resonances likely occur between orbitals that are close in real space. This also provides another explanation for why the tails exist even for high disorder strengths. We will further reinforce this conjecture through calculations of the charge transport capacity in the second half of this work.
III.4.3 LIOM Corrections: Quasiperiodic Model
@@
Figure 6: Analysis of the Frobenius norm of the first-order corrections to the noninteracting LIOMs, $\|\tilde{n}_{\alpha}^{(1)}\|_{F}$, for the weakly-interacting quasiperiodic model. (a)-(d): Distributions of $\ln(\|\tilde{n}_{\alpha}^{(1)}\|_{F})$ based on 500 phase shifts $\delta$, sampled at potential depths $h=0.6,1.6,2.6,3.1$, with $\alpha=1\dots L$. Each curve represents the distribution calculated at different system sizes, from $L=20$ to $L=100$ in steps of 10. The inset in (d) shows the same data as in the main (d) panel but with the vertical scale shown on a log scale. The red line in the inset represents a polynomial $\|\tilde{n}_{\alpha}^{(1)}\|_{F}^{-1.75}$ fit to the tail of the probability distribution of $\|\tilde{n}_{\alpha}^{(1)}\|_{F}$— see the text surrounding Eq. (24) for details on the form plotted. Similarly to what was observed for the interacting Anderson model, the quasiperiodic model seems to admit an extensive number of quasiconservation laws for the localized phase $h>2$. (e): Median $\mathrm{med}\|\tilde{n}_{\alpha}^{(1)}\|_{F}$ as a function of $L$ on a log-log scale. Each colored curve represents a different potential depth $h$, from $h=0.1$ to $h=3.1$ in steps of 0.5. One observes that the curves for $h<2$ grow with $L$, indicating that the corrections are, unsurprisingly, delocalized. For $h>2$ the curves saturate, indicating that the noninteracting LIOMs persist to first order when interactions are added to the localized regime of the quasiperiodic model. For a full discussion of the results, see Sec. III.4.3
@@
We note that the $\mathrm{med}\|\tilde{n}^{(1)}_{\alpha}\|_{F}$ curves exhibit a clear finite-size scaling collapse, analogous to the behavior observed in the Anderson model. For completeness, details of the scaling analysis for the quasiperiodic model across all quantities examined in this work are provided in Appendix K.
@@
III.4.4 LIOM Corrections: Conclusions
Our numerical results show that, even when perturbed by weak interactions, the noninteracting LIOMs continue to exist to first order: specifically, we saw that the two localized models considered still preserve an $O(1)$ fraction of the noninteracting LIOMs up to first order. As previously mentioned, these LIOMs hence survive until time scales of $O(\lambda^{-4})$. From the perspective of integrability breaking, this makes interactions a “weak” perturbation to localized models– consistent with the existence of MBL. However, the results raise many questions. In particular, the slow tails in the distributions of $\|\tilde{n}^{(1)}_{\alpha}\|_{F}$ show that perturbation theory breaks down for a finite fraction of the LIOMs— the fate of those LIOMs remains unclear when interactions are added. From just the norms of the corrections, it is unclear what the spatial structures of the resonances are that cause the breakdown of these LIOMs—in particular, do the resonances mix LIOMs locally, or nonlocally? In principle, some such locality information is contained in the operator content of the corresponding corrections, but it is not easy to analyze. In the next section, we introduce the charge transport capacity, which can frame this question from a different and very physically relevant perspective.
@@ IV.1 Charge Transport Capacity
The maximal net charge transferred across $i_{0}$ over time $t$ (optimized over all possible initial states) is the operator norm of
$\Delta\hat{N}_{i_{0}}(t)\equiv\hat{N}_{R}(i_{0},t)-\hat{N}_{R}(i_{0},0)=\int_{0}^{t}\hat{J}_{i_{0}}(s)ds,$ (26)
@@
To make the charge transport practically computable and to quantify this intuition for the MBL phase, we note that the relation
@@
To tightly bound the charge transport, ideally, one would want to find $\hat{D}$ such that the operator norm of $\hat{M}_{i_{0}}$ is minimized. In practice, minimizing the operator norm of $\hat{M}_{i_{0}}$ is challenging. Therefore, as we will discuss later, we instead choose to minimize its Frobenius norm, which is much simpler. It does not provide an optimally tight upper bound to $\|\Delta\hat{N}_{i_{0}}(t)\|_{\text{op}}$, but gives a good probe of localization regardless.
[…]
$\big(\hat{M}_{i_{0}}\big)_{\mu\nu}=\frac{(\hat{J}_{i_{0}})_{\mu\nu}}{i(E_{\mu}-E_{\nu})}=\big(\hat{N}_{R}\big)_{\mu\nu}\penalty 10000\ .$ (29)
@@
$\hat{M}_{i_{0}}=[\hat{N}_{R}(i_{0})]_{\text{off-diag}}=\hat{N}_{R}-[\hat{N}_{R}]_{\text{diag}}\penalty 10000\ .$ (30)
@@
From now on, we will consider $\hat{M}_{i_{0}}$ given by Eq. (30) and we will call this operator the charge transport capacity operator, and its operator norm the charge transport capacity, as norm of $\hat{M}$ provides a strict upper bound on the number of particles that can cross $i_{0}$, under any process and at any time. Consequently, if $\|\hat{M}\|_{\text{op}}$ remains finite in the thermodynamic limit, it serves as a clear signature of nonergodicity: regardless of how the system is initialized or how long it evolves, no more than a finite number of particles can ever be transported across $i_{0}$. We also claim that localization (in the sense of the Hamiltonian being described by an l-bit model) implies that $\hat{M}$ is quasilocal around $i_{0}$. This illustrates the power of $\hat{M}$ to describe localization: if its norm is finite in the thermodynamic limit, that would imply nonergodicity. Conversely, localization also implies that its norm would be finite.
@@
We note that the above argument relies on the validity of the LIOM model in the MBL phase. However, the possible double-logarithmic growth of the number entropy in the MBL regime [11, 79, 80, 81] suggests that more subtle transport phenomena may persist even deep within the localized phase; whether this is true is not entirely resolved in the literature [82, 83, 5]. Regardless, we emphasize that our calculations of the charge transport in this paper are fully independent of any assumptions about the form of the LIOM description in the localized regime.
We briefly note that the charge transport capacity introduced here does not have a straightforward relation to the more commonly-studied DC conductivity. For example, scaling analyses of conductivity in 1D localized systems were carried out in Ref. [84]. We suspect that the two quantities capture related aspects of transport and may share a deeper connection, but we refrain from making any rigorous claims here.
Finally, we note that this operator is easy to calculate numerically if one knows the full spectrum of $\hat{H}$; this could be an interesting new measure to explore in the full interacting problem, although in this work we will primarily focus on perturbative characterizations.
@@ IV.2 Analytical Forms of Noninteracting Transport and Its Perturbative Correction
We thus find that $\hat{M}^{(1)}$ has to satisfy the relation $\hat{H}_{0},\hat{M}^{(1)}]+[\hat{V},\hat{M}^{(0)}]=0$. In this case, $\hat{M}^{(1)}$ is a four-fermion operator:
@@ IV.3.1 Transport in the Single-Particle Anderson Model
Figure 10: Analysis of $\|\hat{M}^{(0)}\|_{F}$ for the noninteracting Anderson Model based on 10000 disorder realizations, open boundary conditions, and with $i_{0}=\lfloor L/2\rfloor$. (a)–(d): Calculated probability density functions of $\|\hat{M}^{(0)}\|_{F}$. Each panel represents different disorder strengths $W=0.6,1.6,2.6,3.1$. Within each panel, each colored curve represents different system sizes from $L=20$ to $L=100$ in steps of $10$. The distributions are converged at large $L$, consistent with the presence of localization for arbitrarily weak disorder. (e): Medians of the distributions in (a)-(d), $\mathrm{med}\|\hat{M}^{(0)}\|_{F}$, as a function of $L$. Each curve represents a different disorder strength. Here, one sees that the curves saturate at large $L$, reflecting that, in the thermodynamic limit, charge cannot freely flow in the single-particle Anderson model due to localization. In the no disorder case $W=0$, the median scales as $\approx\frac{1}{4}\sqrt{L}$.
@@ IV.3.2 Transport in the Single-Particle Quasiperiodic Model
Figure 12: Analysis of the Frobenius norm of the charge transport capacity, $\|\hat{M}^{(0)}\|_{F}$, for the noninteracting quasiperiodic model with open boundary conditions. The analysis is done based on 10000 evenly spaced phase shifts $\delta$ between $\delta=0$ and $\delta=\frac{1}{k}$, with $i_{0}=\lfloor L/2\rfloor$, and wavenumber $k=\sqrt{2}$. (a)–(d): Calculated probability density functions of $\|\hat{M}^{(0)}\|_{F}$. Within each panel, each colored curve represents a different system size, going from $L=20$ to $L=100$, in steps of $10$. Panels (a) and (b) are in the delocalized phase for $h=0.6$ and $1.6$, where one sees narrow and rightwards shifting peaks. This is a signature of unbounded transport in the delocalized phase. The localized phase is represented in panels (c) and (d) for potential depths $h=2.6,3.1$, which shows distributions converging sharply with increasing $L$. This reflects that transport is finite and limited in the localized phase. (e): We show the median of $\|\hat{M}^{(0)}\|_{F}$, $\mathrm{med}\|\hat{M}^{(0)}\|_{F}$, as a function of $L$. Each differently colored curve represents a different potential depth $h$, from $h=0.1$ to $h=3.1$ in steps of 0.5. We also show the curve at the transition $h=2$ (yellow). The curves in the extended phase $h<2$ show a square root scaling similar to that in the no-disorder case, while the curves in the localized phase for $h>2$ saturate to finite values at large $L$.
@@ IV.4.1 Anderson Model
In Fig. 13, we show numerical results for $\|\hat{M}^{(1)}\|_{F}$ in the weakly-interacting Anderson model. Here, we omit results for $W=0.1$ for visual clarity in the plots.
@@
Figure 13: Analysis of the Frobenius norm of the first-order correction to the charge transport capacity, $\|\hat{M}^{(1)}\|_{F}$, for the weakly-interacting Anderson Model across the link with $i_{0}=\lfloor L/2\rfloor$ and open boundary conditions. (a)–(d): Calculated probability density functions of $\ln(\|\hat{M}^{(1)}\|_{F})$ based on 1000 disorder realizations. Each panel shows a different disorder strength $W=0.6,1.6,2.6,3.1$. Within each panel, each colored curve represents a different system size, with system sizes from $L=20$ to $L=100$ in steps of 10. The inset in (d) shows a distribution of $\|\hat{M}^{(1)}\|_{F}$ at $h=3.1$ with a natural log scale on the y and x-axes, with a $\|\hat{M}^{(1)}\|_{F}^{-2}$ (red line) comparison to the tail— see the text surrounding Eq. (24) for details on the form plotted. The convergence of the distributions at large disorder and $L$ indicates that the corrections to transport are bounded in the thermodynamic limit, suggesting that charge transport is suppressed in the presence of interactions at first-order in perturbation theory. (e): Plot of $\mathrm{med}\|\hat{M}^{(1)}\|_{F}$ as a function of $L$. Each curve represents a different disorder strength $W$, from $W=0.6$ to $3.1$ in steps of 0.5. The inset in (e) shows the same data but plotted using a log-log scale. At large $L$, the transport is bounded, indicating the presence of bottlenecks even when interactions are added.
@@
To check this, we numerically calculate whether or not transport corrections larger than the median at a site $i_{0}$ will continue to stay larger than the median as one travels away from $i_{0}$. Thus, we define $\Theta_{i}\equiv\Theta\left(\|\hat{M}_{i}^{(1)}\|_{F}-\mathrm{med}\|\hat{M}_{i}^{(1)}\|_{F}\right)$. Here, $\Theta$ refers to the Heaviside function—thus, $\Theta_{i}$ is 1 if $\|\hat{M}_{i}^{(1)}\|_{F}$ is larger than the median of the $\|\hat{M}_{i}^{(1)}\|_{F}$ distribution, and 0 otherwise. We calculate the correlation coefficient $G(D)$ between $\Theta_{L/2}$ and $\Theta_{L/2+D}$ (for concision, we denote $\lfloor L/2\rfloor=L/2$, which is true in this case since we only perform calculations on lattices of even sizes), and limit ourselves to considering cases, out of $500$ disorder realizations, where $\|\hat{M}_{L/2}^{(1)}\|_{F}$ are large charge transport corrections (defined as corrections greater than the median of $\|\hat{M}_{L/2}^{(1)}\|_{F}$ for those 500 disorder realizations) at the central link $i_{0}=\lfloor L/2\rfloor$. The full expression for this correlation function is